how to find velocity at half of maximum height provided velocity is v and range is max

Grade:11

1 Answers

Vikas TU
7 years ago
Maximum Height = v^2/2g
Half of Max. Height = H/2 =>H’ = v^2/4g
Range is maximum means thetha of projection = 45 degree.
Therfore at half of max. height = > H = u^2/8g
u^2 = 8gH
u = root(8gH) m/s